OpenStudy (anonymous):

f +2/9=9 1/9 What does f equal? A F=8 8/9 B F=7 8/9 C F=8 7/9

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Use the subtraction property of equality and subtract 2/9 from both sides of the equation.

OpenStudy (skullpatrol):

$$\Huge \text{If } a=b,$$$$\Huge \text{then } a-c=b-c.$$

OpenStudy (skullpatrol):

$$\Large f +\frac{2}{9}=9 \frac{1}{9}$$$$\Large f +\frac{2}{9}=9 +\frac{1}{9}$$$$\Large f =9 +\frac{1}{9}-\frac{2}{9}$$

OpenStudy (skullpatrol):

$$\Large f +\frac{2}{9}=9 \frac{1}{9}$$ $$\Large f +\frac{2}{9}= \frac{82}{9}$$ $$\Large f = \frac{82}{9}-\frac{2}{9}$$ $$\Large f=\frac{82-2}{9}$$$$\Large f=\frac{80}{9}$$